OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation Y=5x-2

OpenStudy (anonymous):

A. –2 B. 5 C. –5 D. - 1/5

OpenStudy (anonymous):

For two lines to be perpendicular, we know that the slope of line 1 must be equal to the negative reciprocal of the slope of line two. i.e.: \[(L1 \perp L2) \rightarrow (m1 = -\frac{ 1 }{ m2 }) \rightarrow (m2 = -\frac{1}{m1})\] So, to find the slop of the second line, we first need to know the slop of the first line (m1) Lucky for us our equation is already in the form: \[y=mx+b\] We know that in this form the slope is equal to 'm'. In this case m1 = 5. Plug this into our equation from above: \[m2=-\frac{1}{m1}\] \[m2=-\frac{1}{5}\] The slope of a line, perpendicular to \[y=5x-2\] is \[-\frac{1 }{ 5}\] Some Notes: -notice that the '-2' in the original equation doesn't come into play at all. Perpendicular doesn't depend on 'b'. It could have been anything and we would have gotten the same answer. -If we wanted parallel instead of perpendicular it would have been even easier: m1=m2.

OpenStudy (anonymous):

Thank you...that helps alot! (:

OpenStudy (anonymous):

Got time for a few more?

OpenStudy (anonymous):

Ya for sure.

OpenStudy (anonymous):

Sorry bout that my computer is acting up..just give me a sec here(:

OpenStudy (anonymous):

What is the Slpoe of a line paralell to the line with equation 5x-2y=11? A 5/2 B-5/2 C-11/2 D-2/5

OpenStudy (anonymous):

First to make this easier lets put the equation in the form \[y=mx+b\] We do this by solving for y. \[5x-2y=11\]\[2y=5x-11\]\[y=\frac{5}{2}x-\frac{11}{2}\] Good so far?

OpenStudy (anonymous):

Yeah..im following so far(:

OpenStudy (anonymous):

Ok. So with the equation in this form we can see that our slope, m1 is equal to \[\frac{5}{2}\] In order for two lines to be parallel, the must have exactly the same slope. i.e.: \[m_{1}=m_{2}\] We just found out \[m_{1}=\frac{5}{2}\] so the answer is just \[m_{2}=m_{1}=\frac{5}{2}\] The answer is A.

OpenStudy (anonymous):

Ok... Yeah that does make sence:P Isnt Slope like have something to do with rise over run or is that something else?

OpenStudy (anonymous):

The slope of any strait line is rise over run. As a formula: \[m=\frac{ rise }{ run }=\frac{ \Delta y }{ \Delta x}=\frac{ y _{2}-y_{1} }{ x_{2}-x_{1} }\]

OpenStudy (anonymous):

alright i think i can remember that...maybe. Which of the following has a slope of -2 and a Y- intercept of 4? A.Y=2x-4 B.Y=-2x-4 C. Y=2x+4 D.Y=-2x+4

OpenStudy (anonymous):

So all the equations are already in the form \[y=mx+b\] In this formula we know that \[m=slope\]and \[b=Yintercept\] So we can just plug in the values they give us: \[y=mx+b\]\[y=−2x+4\] The answer is 'D'.

OpenStudy (anonymous):

Ok i think i understand all that..one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Which ordered pair is on teh praph of the equation 2x+5y=42 A(0,1) B. (7,-2) C. (3,-2) D. (0,-1)

OpenStudy (anonymous):

*the graph* sorry

OpenStudy (anonymous):

The easiest way is to just just check them all one at a time. Let go: A. \[2x+5y=42\]\[2(0)+5(1)=42\]\[2+5=42\]\[7=42\] Nope.

OpenStudy (anonymous):

Ok hold on.....none of them are going to work. Are you sure the question is right?

OpenStudy (anonymous):

let me check again...

OpenStudy (anonymous):

Yeah that is what the question was no wonder i got it wrong...sorry for the inconvenience. but thanks for the help with the other problems.

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Thanks Bye(: